### v2

# An improved perturbative QCD study of the decays $B_{c}^{+}\to\eta_{c}L^{+}$

Abstract
We perform an improved perturbative QCD study of the decays $B_{c}^{+}\to\eta_{c}L^{+}$ ($L$ denotes the light ground state pseudoscalar, vector mesons and the corresponding $p$-wave scalar, axial-vector, and tensor ones) and predict their branching ratios (BRs) associated with relative ratios at leading order in the strong coupling $\alpha_{s}$. Our results ${\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})=(2.03^{+0.53}_{-0.41})\times 10^{-3}$ and ${\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})/{\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})=1.74^{+0.66}_{-0.50}$ are consistent with several available predictions in different approaches within uncertainties. Inputting the measured $\eta_{c}\to p\bar{p}$ and $\eta_{c}\to\pi^{+}\pi^{-}(\pi^{+}\pi^{-},K^{+}K^{-},p\bar{p})$ BRs with $p$ here being a proton, we derive the multibody $B_{c}^{+}\to\eta_{c}(\pi,\rho)^{+}$ BRs through secondary decay chains via resonance $\eta_{c}$ under the narrow-width approximation, which might facilitate the (near) future tests of $B_{c}\to\eta_{c}$ decays. Under the $q\bar{q}$ assignment for light scalars, different to $B_{c}$ decaying into $J/\psi$ plus a scalar meson and other $B_{c}^{+}\to\eta_{c}L^{+}$ modes, surprisingly small $\Delta S=0$ BRs around ${\cal O}(10^{-7}-10^{-9})$ and highly large ratios near ${\cal O}(10^{2})$ between the $\Delta S=1$ and $\Delta S=0$ BRs are found in the $B_{c}$ decays to $\eta_{c}$ plus light scalars, with $S$ being strange number. Many large BRs and interesting ratios presented in this work could be tested by the Large Hadron Collider experiments, which would help us to examine the reliability of this improved perturbative QCD formalism for $B_{c}$-meson decays and further understand the QCD dynamics in the considered decay modes, as well as in the related hadrons.
pacs:
13.25.Hw, 12.38.Bx, 14.40.Nd
preprint: JSNU-PHY-HEP-OCT.25
I Introduction
It is well known that, disparate from the $b\bar{b}$ and $c\bar{c}$ states, the $B_{c}$ meson is flavor-antisymmetric while unique since it is the only ground state containing two different heavy quarks $b$ and $c$ simultaneously Brambilla:2004wf ; Brambilla:2010cs . Peculiar to the more extensively studied $B_{u}$, $B_{d}$, and $B_{s}$ mesons, both constituents in a $B_{c}$ meson can decay separately, which offers a precious opportunity to expand our understanding of heavy $B$-meson physics through thoroughly exploring the more rich and complicated quantum chromodynamics (QCD) in the perturbative and nonperturbative regimes. Thus, the weak decays of the $B_{c}$ meson has prompted significant theoretical and experimental interest since its first discovery at Tevatron in 1998 CDF:1998axz ; CDF:1998ihx .
On the experimental side, the Large Hadron Collider (LHC) experiments starting running in 2009 have measured many $B_{c}$-meson decay channels of interest, such as $B_{c}^{+}\to J/\psi(\pi,2\pi,3\pi)$, even $B_{c}^{+}\to\chi_{cJ}\pi^{+}(J=0,1,2)$, etc. ParticleDataGroup:2024cfk ; HeavyFlavorAveragingGroupHFLAV:2024ctg . The $\eta_{c}$ meson, the lowest-lying $c\bar{c}$ pseudoscalar state, has attracted considerable theoretical and experimental attention since its discovery Partridge:1980vk . It decays primarily via $c\bar{c}$ annihilation into two gluons and is expected to have numerous hadronic decay modes into two- or three-body hadronic charged and/or neutral final states, which, unfortunately, seems not friendly for experimental studies at LHC experiments. Hence, though struggling against the large background and the small efficiency at LHC for these problems, the $B_{c}^{+}\to\eta_{c}\pi^{+}$ decay is not yet observed presently. However, along with the successful upgrade and resumed running of Large Hadron Collier-beauty (LHCb) detector since 2022, the ensuingly exciting results are forthcoming Gao:2024lhc . The continuously collected data with this accomplished even the future prospective upgrades of LHCb detector certainly will offer us a precious opportunity to promote the $B_{c}$-meson physics into a precision era. It is therefore expected that a huge amount of data with superior quality can facilitate a promising measurement of $B_{c}^{+}\to\eta_{c}\pi^{+}$, for example, via $\eta_{c}\to p\bar{p}$ decay chain. Here, $p$ denotes a proton.
[…]
Table 1: Various predictions for ${\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})$, ${\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})/{\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})$ and ${\rm BR}(B_{c}^{+}\to\eta_{c}K^{+})/{\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})$ in the literature.
Observables Chang:1992pt Anisimov:1998uk Colangelo:1999zn AbdElHady:1999xh Verma:2001hb Ebert:2003cn Ivanov:2006ni Hernandez:2006gt Choi:2009ym  [^1] The results are calculated on the basis of Coulomb plus linear confining and harmonic oscillator (in the parentheses) potentials. Naimuddin:2012dy Qiao:2012hp  [^2] These results are calculated with the non-relativistic QCD approach at leading order in the strong coupling $\alpha_{s}$. And the corresponding next-to-leading order BR in $\alpha_{s}^{2}$ is $(5.19^{+0.70}_{-1.07})\times 10^{-3}$. Rui:2014tpa Issadykov:2018myx Nayak:2022qaq Wu:2024gcq Deng:2025znr  [^3] These results are calculated with the QCD factorization approach in $\alpha_{s}$. And the corresponding next-to-next-to-leading order BR in $\alpha_{s}^{2}$ is $(0.81^{+0.88}_{-0.56})\times 10^{-3}$.
$10^{3}\cdot{\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})$ $1.80$ $1.30$ $0.26$ $1.40$ $9.30$ $0.85$ $1.90$ $0.94$ $0.91(1.16)$ $0.34$ $2.95$ $2.98^{+1.24}_{-1.05}$ $1.89^{+0.37}_{-0.37}$ $0.397$ $1.40^{+0.40}_{-0.40}$ $0.79^{+0.86}_{-0.55}$
$\frac{{\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})}{{\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})}$ $1.00$ $1.78$ $0.19$ $1.27$ $2.07$ $1.39$ $1.12$ $1.24$ $\dots$ $1.00$ $1.33$ $1.28^{+0.69}_{-0.56}$ $1.87^{+0.52}_{-0.52}$ $1.02$ $0.56^{+0.21}_{-0.21}$ $0.95^{+1.04}_{-0.68}$
$10^{2}\cdot\frac{{\rm BR}(B_{c}^{+}\to\eta_{c}K^{+})}{{\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})}$ $7.78$ $10.00$ $8.00$ $7.86$ $5.05$ $8.24$ $7.89$ $7.98$ $8.13(8.10)$ $8.82$ $7.12$ $8.05^{+0.68}_{-0.99}$ $7.94^{+0.02}_{-0.05}$ $7.81$ $7.86^{+0.14}_{-0.08}$ $7.51^{+0.02}_{-0.00}$
Table 2: Same as Table 1 but for $B_{c}^{+}\to\eta_{c}(\rho,K^{*})^{+}$.
Observables Chang:1992pt Anisimov:1998uk Colangelo:1999zn AbdElHady:1999xh Verma:2001hb Ebert:2003cn Ivanov:2006ni Hernandez:2006gt Choi:2009ym Naimuddin:2012dy Qiao:2012hp Rui:2014tpa Issadykov:2018myx Nayak:2022qaq Wu:2024gcq Deng:2025znr
$10^{3}\cdot{\rm BR}(B_{c}^{+}\to\eta_{c}\rho^{+})$ $4.90$ $3.00$ $0.67$ $3.30$ $3.70$ $2.10$ $4.50$ $2.40$ $2.57(3.24)$ $1.06$ $7.89$ $9.83^{+3.33}_{-2.59}$ $5.18^{+1.04}_{-1.04}$ $1.24$ $3.80^{+0.10}_{-0.10}$ $2.15^{+2.23}_{-1.46}$
$\frac{{\rm BR}(B_{c}^{+}\to\eta_{c}\rho^{+})}{{\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})}$ $2.72$ $4.11$ $0.52$ $3.0$ $82.2$ $3.44$ $2.65$ $3.16$ $\dots$ $3.12$ $3.55$ $4.22^{+2.05}_{-1.57}$ $5.13^{+1.45}_{-1.45}$ $3.18$ $1.52^{+0.37}_{-0.37}$ $2.60^{+2.71}_{-1.80}$
$10^{2}\cdot\frac{{\rm BR}(B_{c}^{+}\to\eta_{c}K^{*+})}{{\rm BR}(B_{c}^{+}\to\eta_{c}\rho^{+})}$ $5.10$ $7.00$ $5.97$ $5.45$ $3.70$ $5.24$ $5.56$ $5.42$ $5.06(5.25)$ $5.66$ $5.20$ $5.80^{+0.52}_{-0.45}$ $5.60^{+0.03}_{-0.04}$ $5.24$ $5.53^{+1.14}_{-1.48}$ $5.12^{+0.15}_{-0.06}$
To the best of our knowledge, the nature of light scalars, especially those under or near 1 GeV, remains a long-standing puzzle in hadron physics. Recently, the CMS Collaboration found strong evidence of $f_{0}(980)$ being a normal quark-antiquark state CMS:2023rev , even though the ALICE Collaboration supported the $K_{0}^{*}(700)$ being a four-quark state ALICE:2023eyl . Undoubtedly, more endeavors need to be devoted to this field. The useful clues about the nature of light scalars could be collected indirectly through probing their productions in the heavy hadron decays, just like the $B\to f_{0}(980)K$ decays observed in the $B$-factory experiments Abe:2002av ; Aubert:2003mi . So far, there are two different scenarios to describe these scalar mesons $a_{0}(980,1450)$ and $K_{0}^{*}(700,1430)$ under the $q\bar{q}$ assignment Cheng:2005nb . In scenario 1 ($S1$), the scalar mesons $a_{0}(980)$ and $K_{0}^{*}(700)$ are treated as the lowest-lying states, and those $a_{0}(1450)$ and $K_{0}^{*}(1430)$ are the first excited states correspondingly. And, in scenario 2 ($S2$), the scalar mesons $a_{0}(1450)$ and $K_{0}^{*}(1430)$ are viewed as the ground states, while those $a_{0}(980)$ and $K_{0}^{*}(700)$ might be the four-quark states. However, as stressed in Cheng:2005nb , it is difficult in practice to make quantitative predictions based on the four-quark or tetraquark picture for light scalars because the calculations of decay constant and form factors of light scalars are beyond the conventional quark model and the involved nonfactorizable contributions cannot be calculated in the available QCD-based factorization framework. Moreover, the productions of light scalars from the vacuum in the $B_{c}^{+}\to\eta_{c}S^{+}$ decays are expected to be highly suppressed originating from the nearly-zero vector decay constants $f_{S}$ ( Actually, the vector decay constants $f_{S}=0$ in the $SU(3)$ limit.) Cheng:2005nb . It means that investigations on these decays must go beyond naive factorization. Therefore, the predictions in this work are made on the basis of two-quark model for light scalars within the iPQCD framework.
The $p$-wave light axial-vectors have been investigated at both experimental and theoretical aspects. However, our understanding about their nature is still far from complete Du:2022nno . In the spectroscopy study ParticleDataGroup:2024cfk , $a_{1}(1260)$ and $b_{1}(1235)$ are the $1^{3}\!P_{1}$ and $1^{1}\!P_{1}$ axial-vector states, respectively, carrying quantum numbers $J^{PC}=1^{++}$ and $J^{PC}=1^{+-}$ correspondingly. While it is very interesting that the strange $K_{1}(1270)$ and $K_{1}(1400)$ mesons are generally regarded as the mixtures of $1^{3}\!P_{1}$ state $K_{1A}$ and $1^{1}\!P_{1}$ state $K_{1B}$ due to the $SU(3)$ flavor broken symmetry Cheng:2011pb ,
$\left(\begin{array}[]{c}|K_{1}(1270)\rangle\\ |K_{1}(1400)\rangle\\ \end{array}\right)$ $=$ $\left(\begin{array}[]{cc}\sin{\theta_{K}}&\hskip 7.96674pt\cos{\theta_{K}}\\ \cos{\theta_{K}}&-\sin\theta_{K}\end{array}\right)\left(\begin{array}[]{c}|K_{1A}\rangle\\ |K_{1B}\rangle\\ \end{array}\right)\;,$ (7)
with mixing angle $\theta_{K}$. The value of $\theta_{K}$ can be related to the masses of the $K_{1}(1270)$ and $K_{1}(1400)$, to the strong decays of the $K_{1}(1270)$ and $K_{1}(1400)$, and to rates of weak decays to final states involving the $K_{1}(1270)$ and $K_{1}(1400)$ Suzuki:1993yc ; Blundell:1995au . Thus, the decays such as $B_{c}^{+}\to(c\bar{c})K_{1}(1270,1400)^{+}$ would be of great interest for exploring the information of $\theta_{K}$. However, presently, there is no consensus on the value of the mixing angle $\theta_{K}$, and the results from various approaches are still quite controversial, e.g., see a short overview in Liu:2024lph ( and references therein). We therefore take both referenced values, i.e., $\theta_{K_{1}}\approx 33^{\circ}$ and $58^{\circ}$ Cheng:2011pb ; Shi:2023kiy into account in the related numerical calculations of this work.
By incorporating the finite charm quark mass effects into Sudakov resummation of the large logarithmic corrections to wave functions through $k_{T}$ resummation at the next-to-leading-logarithm accuracy Liu:2020upy , besides including them in the hard kernel, the iPQCD formalism is now self-consistent for systematically studying the $B_{c}$-meson decays and $B$-meson decaying into charmonia. So far, facilitated by the newly proposed transverse-momentum-dependent $B_{c}$-meson wave function Liu:2018kuo , we have studied the $B_{c}^{+}\to J/\psi M^{+}$ Liu:2023kxr and $B_{c}^{+}\to\chi_{cJ}(P,V)^{+}$ Liu:2025ipe decays, in which, the BRs predicted in the iPQCD formalism are generally consistent with several available predictions in other approaches. Furthermore, the resultant relative ratios such as ${\rm BR}(B_{c}^{+}\to J/\psi a_{1}(1260)^{+}(\to\pi^{+}\pi^{-}\pi^{+}))/{\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})$, ${\rm BR}(B_{c}^{+}\to J/\psi\rho^{+}(\to\pi^{+}\pi^{0}))/{\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})$, ${\rm BR}(B_{c}^{+}\to\chi_{c2}\pi^{+})/{\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})$ etc. agree well with the current data reported by the LHC experiments within theoretical uncertainties. In light of these successful output, we shall analyze the decays $B_{c}^{+}\to\eta_{c}L^{+}$ in the iPQCD formalism at leading order in the strong coupling $\alpha_{s}$ Liu:2018kuo ; Liu:2020upy . Inputting the BRs ParticleDataGroup:2024cfk of strong decays $\eta_{c}\to p\bar{p}$ and $\eta_{c}\to\pi^{+}\pi^{-}(\pi^{+}\pi^{-},K^{+}K^{-},p\bar{p})$, we additionally present the BRs of multibody modes arising from $B_{c}^{+}\to\eta_{c}(\pi,\rho)^{+}$ via resonance $\eta_{c}$ under the narrow-width approximation. The (near) future tests of our iPQCD predictions at experimental facilities are expected to help us to further understand the QCD dynamics involved in these $B_{c}$-meson decays, even explore the inner structure of $\eta_{c}$ and related light hadrons.
The rest of this paper is organized as follows. In Sect. II, the formalism and the perturbative calculations in association with factorization formulas of $B_{c}^{+}\to\eta_{c}L^{+}$ are presented. The numerical results and phenomenological analyses are given in Sect. III. Sect. IV summarizes our main conclusions.
II Formalism and perturbative calculations
For the $B_{c}^{+}\to\eta_{c}L^{+}$ decays, the related weak effective Hamiltonian $H_{{\rm eff}}$ at the quark level can be written as Buchalla:1995vs
$H_{\rm eff}\,$ $=$ $\,{G_{F}\over\sqrt{2}}\biggl\{V^{*}_{cb}V_{uq}[C_{1}(\mu)O_{1}(\mu)+C_{2}(\mu)O_{2}(\mu)]\biggr\}+{\rm H.c.}\;,$ (8)
with the Fermi constant $G_{F}=1.16639\times 10^{-5}{\rm GeV}^{-2}$, the Cabibbo-Kobayashi-Maskawa (CKM) matrix elements $V$, and the Wilson coefficients $C_{i}(\mu)$ at the renormalization scale $\mu$. The local four-quark tree operators $O_{1}$ and $O_{2}$ are read as
$O_{1}\,$ $=$ $\,\bar{q}_{\alpha}\gamma_{\mu}(1-\gamma_{5})u_{\beta}\;\bar{c}_{\beta}\gamma_{\mu}(1-\gamma_{5})b_{\alpha}\;,\qquad O_{2}\,=\,\bar{q}_{\alpha}\gamma_{\mu}(1-\gamma_{5})u_{\alpha}\;\bar{c}_{\beta}\gamma_{\mu}(1-\gamma_{5})b_{\beta}\;,$ (9)
where $q$ denotes the light down quark $d(s)$ for the CKM-favored (-suppressed) processes.
Similar to $B_{c}^{+}\to J/\psi M^{+}$ decays Liu:2023kxr , the kinematics of $B_{c}^{+}\to\eta_{c}L^{+}$ could be defined in the light-cone coordinates as
$P_{1}$ $=$ $\frac{m_{B_{c}}}{\sqrt{2}}(1,1,{\bf 0}_{T})\;,\qquad P_{2}=\frac{m_{B_{c}}}{\sqrt{2}}(1-r_{3}^{2},r_{2}^{2},{\bf 0}_{T})\;,\qquad P_{3}=\frac{m_{B_{c}}}{\sqrt{2}}(r_{3}^{2},1-r_{2}^{2},{\bf 0}_{T})\;.$ (10)
where the ratios $r_{2}=m_{L}/m_{B_{c}}$ and $r_{3}=m_{\eta_{c}}/m_{B_{c}}$ with $m_{L}(m_{\eta_{c}})$ being the light ($\eta_{c}$) meson mass, $P_{1}$ denotes the momentum carried by $B_{c}$ meson in its rest frame, and $P_{2}$ and $P_{3}$ denotes the momenta carried by $L$ and $\eta_{c}$ mesons moving along the plus and minus $z$-directions, respectively. Thanks to conservation of the angular momentum, the possible polarization vectors $\epsilon_{2L}$ ( Here, the subscript $L$ stands for the longitudinal polarization. Not to be confused with the abbreviation $L$ of light mesons.) could be easily derived through the constraints $P_{2}\cdot\epsilon_{2L}=0$ and $\epsilon_{2L}^{2}=-1$. Notice that, if $L$ is a tensor meson, then a new longitudinal polarization vector could be constructed similarly relative to those of vectors and axial-vectors but with an additional factor $\sqrt{2/3}$ Berger:2000wt ; Datta:2007yk ; Liu:2017cwl ; Liu:2023kxr . Then, specifically, $\epsilon_{2L}=\frac{1}{\sqrt{2(1-r_{3}^{2})}r_{2}}(1-r_{3}^{2},-r_{2}^{2},{\bf 0}_{T})$ for vectors and axial-vectors, while $\epsilon_{2L}=\frac{1}{\sqrt{3(1-r_{3}^{2})}r_{2}}(1-r_{3}^{2},-r_{2}^{2},{\bf 0}_{T})$ for tensors in this work. The momenta of valence quarks in the initial- and final-state mesons are parameterized as
$k_{1}$ $=$ $(x_{1}P_{1}^{+},x_{1}P_{1}^{-},{\bf k}_{1T})\;,\qquad k_{2}=(x_{2}P_{2}^{+},x_{2}P_{2}^{-},{\bf k}_{2T})\;,\qquad k_{3}=(x_{3}P_{3}^{+},x_{3}P_{3}^{-},{\bf k}_{3T})\;,$ (11)
where $x_{i}(i=1,2,3)$ is the corresponding momentum fraction.
The $B_{c}^{+}\to\eta_{c}L^{+}$ decay amplitude in the iPQCD formalism can therefore be conceptually written as follows,
$A(B_{c}^{+}\to\eta_{c}L^{+})$ $\sim$ $\int\!\!dx_{1}dx_{2}dx_{3}b_{1}db_{1}b_{2}db_{2}b_{3}db_{3}$ (12)
$\cdot{\mathrm{Tr}}\left[C(t)\Phi_{B_{c}}(x_{1},b_{1})\Phi_{L}(x_{2},b_{2})\Phi_{\eta_{c}}(x_{3},b_{3})H(x_{i},b_{i},t)e^{-S(t)}\right]\;,$
where $b_{i}$ is the conjugate space coordinate of transverse momentum $k_{iT}$; $t$ is the largest running energy scale in hard kernel $H(x_{i},b_{i},t)$; Tr denotes the trace over Dirac and $SU(3)$ color indices; $C(t)$ stands for the Wilson coefficients including the large logarithms $\ln(m_{W}/t)$ Keum:2000ph ; and $\Phi$ is the wave function describing the hadronization of quark and anti-quark to the meson. The Sudakov factor $e^{-S(t)}$ arises from $k_{T}$ resummation, which provides a strong suppression on the long distance contributions in the small $k_{T}$ (or large $b$) region Botts:1989kf . The detailed discussions for $e^{-S(t)}$ can be easily found in the original Refs. Botts:1989kf ; Liu:2018kuo ; Liu:2020upy . Thus, with Eq. (12), we can give the convoluted amplitudes of the decays $B_{c}^{+}\to\eta_{c}L^{+}$ explicitly through the evaluations of hard kernel $H(x_{i},b_{i},t)$ at leading order in the $\alpha_{s}$ expansion within the iPQCD formalism.
The wave function for $B_{c}$ meson with a “heavy-light” structure can generally be defined as Liu:2018kuo ; Keum:2000ph ; Lu:2002ny
$\Phi_{B_{c}}(x,{\bf k}_{T})$ $=$ $\frac{i}{\sqrt{2N_{c}}}\biggl\{(P\hskip-7.39772pt/+m_{B_{c}})\gamma_{5}\phi_{B_{c}}(x,{\bf k}_{T})\biggr\}\;,$ (13)
where $P$ is the momentum of $B_{c}$ meson, $N_{c}=3$ is the color factor, and $x$ and ${\bf k}_{T}$ are the momentum fraction and intrinsic transverse momentum of charm quark in the $B_{c}$ meson. Here, $\phi_{B_{c}}(x,{\bf k}_{T})$ is the $B_{c}$-meson leading-twist distribution amplitude, whose explicit form in the impact $\bf b$} space is as the following Liu:2018kuo ,
$\phi_{B_{c}}(x,{\bf b})$ $=$ $\frac{f_{B_{c}}}{2\sqrt{2N_{c}}}N_{B_{c}}x(1-x)\exp\left[-\frac{(1-x)m_{c}^{2}+xm_{b}^{2}}{8\beta_{B_{c}}^{2}x(1-x)}\right]\exp\left[-2\beta_{B_{c}}^{2}x(1-x){\bf b}^{2}\right]\;,$ (14)
with the decay constant $f_{B_{c}}=0.489\pm 0.005$ GeV Chiu:2007km , the shape parameter $\beta_{B_{c}}=1.0\pm 0.1$ GeV Liu:2018kuo , and $m_{c}$ and $m_{b}$ the charm and bottom quark masses. Moreover, the normalization constant $N_{B_{c}}$ is fixed by the following relation,
$\int_{0}^{1}\phi_{B_{c}}(x,{\bf b}=0)dx\equiv\int_{0}^{1}\phi_{B_{c}}(x)dx$ $=$ $\frac{f_{B_{c}}}{2\sqrt{2N_{c}}}\;.$ (15)
For the $\eta_{c}$ meson, its wave function has been studied within the non-relativistic QCD approach Bondar:2004sv and derived as,
$\Phi_{\eta_{c}}(x)$ $=$ $\frac{i}{\sqrt{2N_{c}}}\gamma_{5}\biggl\{P\hskip-7.39772pt/\ \phi_{\eta_{c}}^{v}(x)+m_{\eta_{c}}\phi_{\eta_{c}}^{s}(x)\biggl\}\;,$ (16)
with $P$ and $m$ being the momentum and mass of $\eta_{c}$, and $x$ describing the charm-quark momentum fraction in $\phi^{v}_{\eta_{c}}(x)$ and $\phi^{s}_{\eta_{c}}(x)$ the twist-2 and twist-3 distribution amplitudes,
$\phi_{\eta_{c}}^{v}(x)$ $=$ $9.58\frac{f_{\eta_{c}}}{2\sqrt{2N_{c}}}x(1-x){\cal C}(x)\;,\qquad\phi_{\eta_{c}}^{s}(x)=1.97\frac{f_{\eta_{c}}}{2\sqrt{2N_{c}}}{\cal C}(x)\;,$ (17)
where $f_{\eta_{c}}=0.387\pm 0.007$ Becirevic:2013bsa is the decay constant and the function ${\cal C}(x)$ reads
${\cal C}(x)$ $=$ $\biggl[(x(1-x))/(1-4x(1-x)(1-v^{2}))\biggl]^{1-v^{2}}\;.$ (18)
with $v^{2}=0.3$ standing for small relativistic corrections to the Coulomb wave functions.
The light-cone wave functions including distribution amplitudes for light pseudoscalars, scalars, vectors, axial-vectors, and tensors calculated in the QCD sum rules up to twist-3 have been collected in Liu:2023kxr (and references therein). For simplicity, their explicit expressions would no longer be presented in this work. The readers could refer to Liu:2023kxr for detail.
Figure 1: (Color online) Leading order Feynman diagrams for the decays $B_{c}^{+}\to\eta_{c}L^{+}$ in the iPQCD formalism.
The leading-order Feynman diagrams for the decays $B_{c}^{+}\to\eta_{c}L^{+}$ in the iPQCD formalism are shown in Fig. 1. Similar to those in Liu:2023kxr , we use $F_{e}$ and $M_{e}$ to describe the factorizable emission and the nonfactorizable emission amplitudes induced by the (VA)(VA)$(V-A)$ operators. The $B_{c}^{+}\to\eta_{c}L^{+}$ decay amplitude can thus be decomposed into
$A(B_{c}^{+}\to\eta_{c}L^{+})$ $=$ $V_{cb}^{*}V_{uq}(F_{e}\cdot f_{L}+M_{e})\;,$ (19)
in which, the related factorization formulas are given explicitly as follows,
  • For $B_{c}^{+}\to\eta_{c}(P,S)^{+}$ decays,
    $F_{e}(P)$ $=$ $-8\pi C_{F}m_{B_{c}}^{4}\int_{0}^{1}dx_{1}dx_{3}\int_{0}^{\infty}b_{1}db_{1}b_{3}db_{3}\phi_{B_{c}}(x_{1},b_{1})(r_{3}^{2}-1)$ (20)
    $\times\biggl\{\left[r_{3}(r_{b}+2x_{3}-2)\phi_{\eta_{c}}^{s}(x_{3})-(2r_{b}+x_{3}-1)\phi_{\eta_{c}}^{v}(x_{3})\right]h_{a}(x_{1},x_{3},b_{1},b_{3})E_{f}(t_{a})$
    $-\left[2r_{3}(1+r_{c}-x_{1})\phi_{\eta_{c}}^{s}(x_{3})+(r^{2}_{3}(x_{1}-1)-r_{c})\phi_{\eta_{c}}^{v}(x_{3})\right]h_{b}(x_{1},x_{3},b_{1},b_{3})E_{f}(t_{b})\biggr\}\;,$
    where the ratios $r_{b}=m_{b}/m_{B_{c}}$ and $r_{c}=m_{c}/m_{B_{c}}$. The hard function $h_{i}(x_{i},b_{i})$ and the evolution function $E_{f}(t_{i})$ could refer to those expressions in Ref. Liu:2023kxr .
    $M_{e}(P)$ $=$ $-\frac{32}{\sqrt{6}}\pi C_{F}m_{B_{c}}^{4}\int_{0}^{1}dx_{1}dx_{2}dx_{3}\int_{0}^{\infty}b_{1}db_{1}b_{2}db_{2}\phi_{B_{c}}(x_{1},b_{1})\phi_{P}^{A}(x_{2})(r_{3}^{2}-1)$ (21)
    $\times\biggl\{\left[r_{3}(x_{3}-x_{1})\phi_{\eta_{c}}^{s}(x_{3})+(x_{1}+x_{2}-1)\phi_{\eta_{c}}^{v}(x_{3})+r_{3}^{2}(x_{1}-x_{2}-2x_{3}+1)\phi_{\eta_{c}}^{v}(x_{3})\right]$
    $\times E_{f}(t_{c})h_{c}(x_{1},x_{2},x_{3},b_{1},b_{2})+\left[(x_{2}+x_{3}-2x_{1}+r_{3}^{2}(x_{3}-x_{2}))\phi_{\eta_{c}}^{v}(x_{3})\right$.
    $\left.+r_{3}(x_{1}-x_{3})\phi_{\eta_{c}}^{s}(x_{3})\right]h_{d}(x_{1},x_{2},x_{3},b_{1},b_{2})E_{f}(t_{d})\biggr\}\;,$
    and
    $F_{e}(S)$ $=$ $-F_{e}(P)\;,\qquad M_{e}(S)=-M_{e}(P)\;.$ (22)
    but with the corresponding replacement of $\phi_{P}^{A}(x)\to\phi_{S}(x)$ in Eq. (22).
  • For $B_{c}^{+}\to\eta_{c}(V,A,T)^{+}$ decays,
    $F_{e}(V)$ $=$ $8\pi C_{F}m_{B_{c}}^{4}\int_{0}^{1}dx_{1}dx_{3}\int_{0}^{\infty}b_{1}db_{1}b_{3}db_{3}\phi_{B_{c}}(x_{1},b_{1})\sqrt{1-r_{3}^{2}}$ (23)
    $\times\biggl\{\left[r_{3}(r_{b}+2x_{3}-2)\phi_{\eta_{c}}^{s}(x_{3})-(2r_{b}+x_{3}-1)\phi_{\eta_{c}}^{v}(x_{3})\right]h_{a}(x_{1},x_{3},b_{1},b_{3})E_{f}(t_{a})$
    $-\left[2r_{3}(1+r_{c}-x_{1})\phi_{\eta_{c}}^{s}(x3)+(r^{2}_{3}(x_{1}-1)-r_{c})\phi_{\eta_{c}}^{v}(x_{3})\right]h_{b}(x_{1},x_{3},b_{1},b_{3})E_{f}(t_{b})\biggr\}\;,$
    $M_{e}(V)$ $=$ $\frac{32}{\sqrt{6}}\pi C_{F}m_{B_{c}}^{4}\int_{0}^{1}dx_{1}dx_{2}dx_{3}\int_{0}^{\infty}b_{1}db_{1}b_{2}db_{2}\phi_{B_{c}}(x_{1},b_{1})\phi_{V}(x_{2})\sqrt{1-r_{3}^{2}}$ (24)
    $\times\biggl\{\left[(x_{1}+x_{2}-1)\phi_{\eta_{c}}^{v}(x_{3})+r_{3}^{2}(x_{1}-x_{2}-2x_{3}+1)\phi_{\eta_{c}}^{v}(x3)+r_{3}(x_{3}-x_{1})\phi_{\eta_{c}}^{s}(x_{3})\right]$
    $E_{f}(t_{c})h_{c}(x_{1},x_{2},x_{3},b_{1},b_{2})+\left[(x_{2}+x_{3}-2x_{1}+r_{3}^{2}(x_{3}-x_{2}))\phi_{\eta_{c}}^{v}(x_{3})\right$.
    $\left.+r_{3}(x_{1}-x_{3})\phi_{\eta_{c}}^{s}(x_{3})\right]h_{d}(x_{1},x_{2},x_{3},b_{1},b_{2})E_{f}(t_{d})\biggr\}\;,$
    and
    $F_{e}(A)$ $=$ $-F_{e}(V)\;,\qquad F_{e}(T)=0\;,\qquad M_{e}(A)=-M_{e}(V)\;,\qquad M_{e}(T)=\sqrt{\frac{2}{3}}M_{e}(V)\;.$ (25)
    associated with the replacements of $\phi_{V}(x)\to\phi_{A,T}(x)$ in (25) correspondingly.
The corresponding BR is then given, in the rest frame of a heavy $B_{c}$ meson, by
${\rm BR}(B_{c}^{+}\to\eta_{c}L^{+})$ $\equiv$ $\tau_{B_{c}}\cdot\Gamma(B_{c}^{+}\to\eta_{c}L^{+})=\tau_{B_{c}}\cdot\frac{G_{F}^{2}|\bf{P_{c}}|}{16\pi m^{2}_{B_{c}}}|A(B_{c}^{+}\to\eta_{c}L^{+})|^{2}\;,$ (26)
with the $B_{c}$-meson lifetime $\tau_{B_{c}}$ and the decay width $\Gamma$. Note that, $|{\bf P}_{c}|\equiv|{\bf P}_{L}|=|{\bf P}_{\eta_{c}}|=\sqrt{\lambda(m_{B_{c}}^{2},m_{L}^{2},m_{\eta_{c}}^{2})}/(2m_{B_{c}})$ is the momentum of either $L$ or $\eta_{c}$ meson in the final states, with the $\rm K\ddot{a}ll\acute{e}n$} function $\lambda(x,y,z)=x^{2}+y^{2}+z^{2}-2xy-2xz-2yz$ ParticleDataGroup:2024cfk .
III Numerical Results and Discussions
In numerical calculations, central values of the input parameters will be used implicitly unless otherwise stated. We adopt the relevant QCD scale (GeV), masses (GeV), and $B_{c}$-meson lifetime (ps) Keum:2000ph ; ParticleDataGroup:2024cfk
$\Lambda_{\overline{\mathrm{MS}}}^{(f=4)}$ $=$ $0.250\;,\qquad m_{W}=80.41\;,\qquad m_{B_{c}}=6.275\;,\qquad m_{\eta_{c}}=2.98\;,$
$\tau_{B_{c}}$ $=$ $0.507\;,\qquad m_{b}=4.8\;,\qquad m_{c}=1.5\;.$ (27)
and the CKM matrix elements ParticleDataGroup:2024cfk ,
$|V_{cb}|$ $=$ $0.04182^{+0.00085}_{-0.00074}\;,\quad|V_{ud}|=0.97435\pm 0.00016\;,\quad|V_{us}|=0.22500\pm 0.00067\;.$ (28)
In the following context, we will classify the considered decays into two different groups conventionally, that is, factorizable-emission dominated $B_{c}^{+}\to\eta_{c}L^{+}$ and factorizable-emission suppressed $B_{c}^{+}\to\eta_{c}L^{+}$, respectively, to present the related iPQCD predictions and phenomenological insights.
III.1 Factorizable-emission dominated $B_{c}^{+}\to\eta_{c}L^{+}$
Specifically, the factorizable-emission dominated $B_{c}^{+}\to\eta_{c}L^{+}$ decays include $B_{c}^{+}\to\eta_{c}(P,V)^{+}$, $B_{c}^{+}\to\eta_{c}a_{1}(1260)^{+}$, and $B_{c}^{+}\to\eta_{c}K_{1}(1270,1400)^{+}$, respectively. The decays $B_{c}^{+}\to\eta_{c}(P,V)^{+}$ have been studied extensively with various $B_{c}\to\eta_{c}$ form factors in different approaches, however, achieving different individual BRs and the associated relative ratios as given in Tables 1 and 2. It is clear that the values given in the literature have a wide spread. While, to our best knowledge, the $B_{c}^{+}\to\eta_{c}a_{1}(1260)^{+}$ and $\eta_{c}K_{1}(1270,1400)^{+}$ modes have not yet been investigated within the QCD-based factorization framework.
The CP-averaged $B_{c}^{+}\to\eta_{c}(\pi,K)^{+}$ BRs in the iPQCD formalism are presented as,
${\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})$ $=$ $2.03^{+0.51}_{-0.40}(\beta_{B_{c}})^{+0.09}_{-0.08}(f_{M})^{+0.00}_{-0.00}(a_{\pi})^{+0.09}_{-0.07}(V_{cb})\times 10^{-3}\;,$ (29)
${\rm BR}(B_{c}^{+}\to\eta_{c}K^{+})$ $=$ $1.52^{+0.39}_{-0.29}(\beta_{B_{c}})^{+0.07}_{-0.06}(f_{M})^{+0.11}_{-0.10}(a_{K})^{+0.07}_{-0.05}(V_{cb})\times 10^{-4}\;,$ (30)
where the uncertainties are dominated by the shape parameter $\beta_{B_{c}}$ from the $B_{c}$-meson distribution amplitude. Although the $B_{c}^{+}\to\eta_{c}\pi^{+}$ BR is not measured yet, its iPQCD result is consistent generally with several predictions from various models and/or approaches already presented in the literature, for example, see Refs. Chang:1992pt ; Ivanov:2006ni ; Rui:2014tpa ; Issadykov:2018myx ; Wu:2024gcq ; Deng:2025znr . Note that, though having a smaller decay constant $f_{\eta_{c}}\sim 0.387$ than $f_{J/\psi}\sim 0.405$ and the same leading-twist distribution amplitudes in both $\eta_{c}$ and $J/\psi$, the $B_{c}^{+}\to\eta_{c}\pi^{+}$ BR is still larger than the $B_{c}^{+}\to J/\psi\pi^{+}$ one in the iPQCD formalism. In fact, from the numerical results of decay amplitudes shown in Table 3, it is found that ${\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})$ and ${\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})$ are governed by highly different QCD dynamics, that is, the former (latter) determined by the contributions from twist-3 (twist-2) distribution amplitude of $\eta_{c}(J/\psi)$ even with enhancement (reduction) by twist-2 (twist-3) one. Any precise measurements in various experiments to help deeply understand the QCD behavior of these $c\bar{c}$-mesons are urgently demanded.
Table 3: Decay amplitudes (in units of $10^{-3}$GeV-3) of $B_{c}^{+}\to\eta_{c}\pi^{+}$ and $B_{c}^{+}\to J/\psi\pi^{+}$ from different twists in the iPQCD formalism. For simplicity, only the central values are quoted for clarifications.
Modes Decay Amplitudes ($F_{e}$) Decay Amplitudes ($M_{e}$)
$B_{c}^{+}\to\eta_{c}\pi^{+}$ $\phi_{\eta_{c}}^{v}(x)$ $\phi_{\eta_{c}}^{s}(x)$ $\phi_{\eta_{c}}^{v}(x)$ $\phi_{\eta_{c}}^{s}(x)$
$8.43-{\it i}8.56$ $26.42-{\it i}113.08$ $1.00-{\it i}2.61$ $0.01-{\it i}0.01$
$B_{c}^{+}\to J/\psi\pi^{+}$ $\phi_{J/\psi}^{L}(x)$ $\phi_{J/\psi}^{t}(x)$ $\phi_{J/\psi}^{L}(x)$ $\phi_{J/\psi}^{t}(x)$
$30.76+{\it i}101.07$ $12.83-{\it i}9.09$ $1.85+{\it i}4.21$ $0.24-{\it i}0.92$
The ratio between ${\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})$ and ${\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})$ in the iPQCD framework could help measure the mode $B_{c}^{+}\to\eta_{c}\pi^{+}$ experimentally,
$R_{\eta_{c}/{J/\psi}}^{\pi}$ $\equiv$ $\frac{{\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})}{{\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})}=1.74^{+0.66}_{-0.50}\;,$ (31)
Notice that, the results about BR and relative ratio shown in Eqs. (29) and (31) are simultaneously well consistent with those predicted in Issadykov:2018myx within uncertainties. According to the LHCb data LHCb:2014mvo , that is, $\frac{\sigma(B_{c}^{+})}{\sigma(B^{+})}\times{\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})=(6.97\pm 0.13)\times 10^{-6}$ Liu:2025ipe in the fiducial region corresponding to the transverse momentum $0<p_{T}<20$ GeV and the rapidity $2.0<y<4.5$, and the assumption ${\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})\sim{\cal O}(10^{-3})$, then the promising measurements of ${\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})$ in the LHCb experiment might be
$\frac{\sigma(B_{c}^{+})}{\sigma(B^{+})}\times{\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})$ $\sim$ $(1.41^{+0.37}_{-0.29})\times 10^{-5}\;.$ (32)
Based on the data for $\eta_{c}$ strong decays to stable hadrons ParticleDataGroup:2024cfk , one can derive the multibody $B_{c}$-meson decays via $\eta_{c}$ resonance under the narrow-width approximation.
  • (1)
    $\eta_{c}\to p\bar{p}$ and $p\bar{p}\pi^{+}\pi^{-}$,
    The inputs of $\eta_{c}\to p\bar{p}$ and $p\bar{p}\pi^{+}\pi^{-}$ decays with
    ${\cal B}(\eta_{c}\to p\bar{p})$ $=$ $(1.33\pm 0.11)\times 10^{-3}\;,\qquad{\cal B}(\eta_{c}\to p\bar{p}\pi^{+}\pi^{-})=(3.7\pm 0.5)\times 10^{-3}\;,$ (33)
    lead to the three-body $B_{c}^{+}\to\pi^{+}\eta_{c}(\to p\bar{p})$ and five-body $B_{c}^{+}\to\pi^{+}\eta_{c}(\to p\bar{p}\pi^{+}\pi^{-})$ BRs as,
    ${\rm BR}(B_{c}^{+}\to\pi^{+}\eta_{c}(\to p\bar{p}))$ $\equiv$ ${\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})\cdot{\cal B}(\eta_{c}\to p\bar{p})=(2.70^{+0.74}_{-0.59})\times 10^{-6}\;,$ (34)
    ${\rm BR}(B_{c}^{+}\to\pi^{+}\eta_{c}(\to p\bar{p}\pi^{+}\pi^{-}))$ $\equiv$ ${\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})\cdot{\cal B}(\eta_{c}\to p\bar{p}\pi^{+}\pi^{-})=(7.51^{+2.21}_{-1.83})\times 10^{-6}\;,$ (35)
  • (2)
    $\eta_{c}\to 2(\pi^{+}\pi^{-}),2(K^{+}K^{-})$ and $\pi^{+}\pi^{-}K^{+}K^{-}$,
    The inputs of $\eta_{c}\to 2(\pi^{+}\pi^{-})$, $\pi^{+}\pi^{-}K^{+}K^{-}$ and $2(K^{+}K^{-})$ decays with
    ${\cal B}(\eta_{c}\to 2(\pi^{+}\pi^{-}))$ $=$ $(9.6\pm 1.5)\times 10^{-3}\;,\qquad{\cal B}(\eta_{c}\to 2(K^{+}K^{-}))=(1.4\pm 0.4)\times 10^{-3}\;,$
    ${\cal B}(\eta_{c}\to\pi^{+}\pi^{-}K^{+}K^{-})$ $=$ $(8.3\pm 1.8)\times 10^{-3}\;,$ (36)
    result in the five-body $B_{c}^{+}\to\pi^{+}\eta_{c}(\to 2(\pi^{+}\pi^{-}),(\pi^{+}\pi^{-}K^{+}K^{-}),2(K^{+}K^{-}))$ BRs as,
    ${\rm BR}(B_{c}^{+}\to\pi^{+}\eta_{c}(\to\pi^{+}\pi^{-}\pi^{+}\pi^{-}))$ $\equiv$ ${\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})\cdot{\cal B}(\eta_{c}\to\pi^{+}\pi^{-}\pi^{+}\pi^{-})$ (37)
    $=$ $(1.95^{+0.59}_{-0.50})\times 10^{-5}\;,$
    ${\rm BR}(B_{c}^{+}\to\pi^{+}\eta_{c}(\to\pi^{+}\pi^{-}K^{+}K^{-}))$ $\equiv$ ${\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})\cdot{\cal B}(\eta_{c}\to\pi^{+}\pi^{-}K^{+}K^{-})$ (38)
    $=$ $(1.68^{+0.57}_{-0.50})\times 10^{-5}\;,$
    ${\rm BR}(B_{c}^{+}\to\pi^{+}\eta_{c}(\to K^{+}K^{-}K^{+}K^{-}))$ $\equiv$ ${\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})\cdot{\cal B}(\eta_{c}\to K^{+}K^{-}K^{+}K^{-})$ (39)
    $=$ $(2.84^{+1.10}_{-0.99})\times 10^{-6}\;,$
In principle, the related measurements are previously difficult due to the large background and the small efficiency at LHC. However, these BRs around ${\cal O}(10^{-6})$ and above with fully charged final-states are expected to be probed in the near future, since the LHCb detector has accomplished a successful upgrade in 2022.
The ratio between the BRs of $B_{c}^{+}\to\eta_{c}K^{+}$ and $B_{c}^{+}\to\eta_{c}\pi^{+}$ is given theoretically in the iPQCD formalism as,
$R^{\eta_{c}}_{K/\pi}$ $\equiv$ $\frac{{\rm BR}(B_{c}^{+}\to\eta_{c}K^{+})}{{\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})}=(7.49^{+0.54}_{-0.49})\times 10^{-2}\;,$ (40)
which agrees well with the naive anticipation $(7.95\pm 0.04)\times 10^{-2}$ within uncertainties. That is, the decays $B_{c}^{+}\to\eta_{c}(\pi,K)^{+}$ are indeed predominated by the factorizable decay amplitudes. Meanwhile, the $a_{1}^{K}$-term induced $SU(3)$ flavor symmetry-breaking effects arising from leading-twist kaon distribution amplitude in the nonfactorizable decay amplitudes lead to a slight deviation to the naive expectation, as stated in Refs. Liu:2023kxr ; Liu:2025ipe .
The $B_{c}^{+}\to\eta_{c}V^{+}$ BRs in the iPQCD formalism can be read as follows,
${\rm BR}(B_{c}^{+}\to\eta_{c}\rho^{+})$ $=$ $5.37^{+1.42}_{-1.08}(\beta_{B_{c}})^{+0.24}_{-0.24}(f_{M})^{+0.00}_{-0.00}(a_{\rho})^{+0.22}_{-0.19}(V_{cb})\times 10^{-3}\;,$ (41)
${\rm BR}(B_{c}^{+}\to\eta_{c}K^{*+})$ $=$ $3.05^{+0.81}_{-0.62}(\beta_{B_{c}})^{+0.19}_{-0.19}(f_{M})^{+0.02}_{-0.02}(a_{K^{*}})^{+0.12}_{-0.11}(V_{cb})\times 10^{-4}\;,$ (42)
where the theoretical errors are also dominated mainly by the $B_{c}$-meson shape parameter $\beta_{B_{c}}$. These predictions are well consistent with those in Refs. Chang:1992pt ; Ivanov:2006ni ; Issadykov:2018myx ; Deng:2025znr within uncertainties.
By employing Eqs. (33) and (36), the measurable BRs of multibody $B_{c}^{+}\to\eta_{c}\rho^{+}$ decays through resonance state $\eta_{c}$ could be derived under the narrow-width approximation as follows,
${\rm BR}(B_{c}^{+}\to\eta_{c}(\to p\bar{p})\rho^{+})$ $\equiv$ ${\rm BR}(B_{c}^{+}\to\eta_{c}\rho^{+})\cdot{\cal B}(\eta_{c}\to p\bar{p})=(7.14^{+2.03}_{-1.60})\times 10^{-6}\;,$ (43)
${\rm BR}(B_{c}^{+}\to\eta_{c}(\to p\bar{p}\pi^{+}\pi^{-})\rho^{+})$ $\equiv$ ${\rm BR}(B_{c}^{+}\to\eta_{c}\rho^{+})\cdot{\cal B}(\eta_{c}\to p\bar{p}\pi^{+}\pi^{-})=(1.99^{+0.60}_{-0.49})\times 10^{-5}\;,$ (44)
${\rm BR}(B_{c}^{+}\to\eta_{c}(\to\pi^{+}\pi^{-}\pi^{+}\pi^{-})\rho^{+})$ $\equiv$ ${\rm BR}(B_{c}^{+}\to\eta_{c}\rho^{+})\cdot{\cal B}(\eta_{c}\to\pi^{+}\pi^{-}\pi^{+}\pi^{-})$ (45)
$=$ $(5.16^{+1.62}_{-1.34})\times 10^{-5}\;,$
${\rm BR}(B_{c}^{+}\to\eta_{c}(\to\pi^{+}\pi^{-}K^{+}K^{-})\rho^{+})$ $\equiv$ ${\rm BR}(B_{c}^{+}\to\eta_{c}\rho^{+})\cdot{\cal B}(\eta_{c}\to\pi^{+}\pi^{-}K^{+}K^{-})$ (46)
$=$ $(4.46^{+1.55}_{-1.34})\times 10^{-5}\;,$
${\rm BR}(B_{c}^{+}\to\eta_{c}(\to K^{+}K^{-}K^{+}K^{-})\rho^{+})$ $\equiv$ ${\rm BR}(B_{c}^{+}\to\eta_{c}\rho^{+})\cdot{\cal B}(\eta_{c}\to K^{+}K^{-}K^{+}K^{-})$ (47)
$=$ $(7.52^{+2.97}_{-2.66})\times 10^{-6}\;.$
The ratio between the $B_{c}^{+}\to\eta_{c}\rho^{+}$ and $B_{c}^{+}\to(J/\psi,\eta_{c})\pi^{+}$ BRs could be deduced as,
$\frac{{\rm BR}(B_{c}^{+}\to\eta_{c}\rho^{+})}{{\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})}$ $=$ $4.59^{+1.77}_{-1.34}\;,$ (48)
and
$R_{\rho/\pi}^{\eta_{c}}$ $\equiv$ $\frac{{\rm BR}(B_{c}^{+}\to\eta_{c}\rho^{+})}{{\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})}=2.65^{+0.02}_{-0.03}\;,$ (49)
Here, the latter ratio $R_{\rho/\pi}^{\eta_{c}}$ is close to that obtained from the $B_{c}^{+}\to J/\psi(\pi,\rho)^{+}$ decays. Moreover, the ratio between the $B_{c}^{+}\to\eta_{c}K^{*+}$ and $B_{c}^{+}\to\eta_{c}\rho^{+}$ BRs is,
$R^{\eta_{c}}_{K^{*}/\rho}$ $\equiv$ $\frac{{\rm BR}(B_{c}^{+}\to\eta_{c}K^{*+})}{{\rm BR}(B_{c}^{+}\to\eta_{c}\rho^{+})}=(5.68^{+0.11}_{-0.11})\times 10^{-2}\;,$ (50)
which matches well with the value $(5.75\pm 0.03)\times 10^{-2}$ anticipated by naive factorization within errors.
Next, for the $B_{c}^{+}\to\eta_{c}a_{1}(1260)^{+}$ channel, its BR in the iPQCD formalism is read as,
${\rm BR}(B_{c}^{+}\to\eta_{c}a_{1}(1260)^{+})$ $=$ $6.91^{+1.84}_{-1.41}(\beta_{B_{c}})^{+0.66}_{-0.64}(f_{M})^{+0.00}_{-0.00}(a_{a_{1}})^{+0.29}_{-0.24}(V_{cb})\times 10^{-3}\;,$ (51)
associated with the relevant ratios,
$\frac{{\rm BR}(B_{c}^{+}\to\eta_{c}a_{1}(1260)^{+})}{{\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})}$ $=$ $5.91^{+2.33}_{-1.81}\;,$ (52)
and
$R_{a_{1}/\rho}^{\eta_{c}}$ $\equiv$ $\frac{{\rm BR}(B_{c}^{+}\to\eta_{c}a_{1}(1260)^{+})}{{\rm BR}(B_{c}^{+}\to\eta_{c}\rho^{+})}=1.29^{+0.51}_{-0.40}\;,\qquad R_{a_{1}/\pi}^{\eta_{c}}\equiv\frac{{\rm BR}(B_{c}^{+}\to\eta_{c}a_{1}(1260)^{+})}{{\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})}=3.40^{+1.32}_{-1.03}\;.$ (53)
The future precise measurements would provide useful information to help study the QCD dynamics among the states pion, $\rho$ and $a_{1}(1260)$ involving the same quark components.
To provide referenced values for experimental measurements of this mode, the $B_{c}^{+}\to\eta_{c}a_{1}(1260)^{+}(\to\pi^{+}\pi^{-}\pi^{+})$ and $B_{c}^{+}\to\eta_{c}a_{1}(1260)^{+}(\to K^{+}K^{-}\pi^{+})$ BRs could be derived in the iPQCD formalism under the narrow-width approximation via resonance state $a_{1}(1260)$ with the strong decays ${\cal B}(a_{1}(1260)^{+}\to\pi^{+}\pi^{-}\pi^{+})=0.50\pm 0.05$ Aubert:2007kpb and ${\cal B}(a_{1}(1260)^{+}\to K^{+}K^{-}\pi^{+})=0.11\pm 0.02$ Liu:2023kxr as,
${\rm BR}(B_{c}^{+}\to\eta_{c}a_{1}(1260)^{+}(\to\pi^{+}\pi^{-}\pi^{+}))$ $\equiv$ ${\rm BR}(B_{c}^{+}\to\eta_{c}a_{1}(1260)^{+})\cdot{\cal B}(a_{1}^{+}\to\pi^{+}\pi^{-}\pi^{+})$ (54)
$=$ $(3.46^{+1.04}_{-0.86})\times 10^{-3}\;,$
and
${\rm BR}(B_{c}^{+}\to\eta_{c}a_{1}(1260)^{+}(\to K^{+}K^{-}\pi^{+}))$ $\equiv$ ${\rm BR}(B_{c}^{+}\to\eta_{c}a_{1}(1260)^{+})\cdot{\cal B}(a_{1}^{+}\to K^{+}K^{-}\pi^{+})$ (55)
$=$ $(7.60^{+2.57}_{-2.21})\times 10^{-4}\;,$
In principle, these large BRs around ${\cal O}(10^{-3})$ could be accessed easily in the related experiments. Moreover, interestingly, Ref. Wang:2012vna ever provided the $B_{c}^{+}\to\eta_{c}\pi^{+}\pi^{-}\pi^{+}$ BR as $1.854\times 10^{-3}$ with $B_{c}\to\eta_{c}$ form factor calculated by QCD sum rules, associated with the relevant ratios between the BRs of $B_{c}^{+}\to\eta_{c}\pi^{+}\pi^{-}\pi^{+}$ and $B_{c}^{+}\to(J/\psi,\eta_{c})\pi^{+}$ as $1.52$ and $1.36$, respectively. For comparison, these values can also be given in the iPQCD formalism,
$\frac{{\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+}\pi^{-}\pi^{+})}{{\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})}$ $=$ $2.96^{+1.20}_{-0.99}\;,\qquad\frac{{\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+}\pi^{-}\pi^{+})}{{\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})}=1.70^{+0.68}_{-0.55}\;,$ (56)
$\frac{{\rm BR}(B_{c}^{+}\to\eta_{c}K^{+}K^{-}\pi^{+})}{{\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})}$ $=$ $0.65^{+0.28}_{-0.23}\;,\qquad\frac{{\rm BR}(B_{c}^{+}\to\eta_{c}K^{+}K^{-}\pi^{+})}{{\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})}=0.37^{+0.16}_{-0.13}\;.$ (57)
These numerical values could be confronted with the future measurements.
The $B_{c}^{+}\to\eta_{c}K_{1}(1270,1400)^{+}$ BRs predicted in the iPQCD formalism with different $\theta_{K}$ are as follows,
${\rm BR}(B_{c}^{+}\to\eta_{c}K_{1}(1270)^{+})$ $=$ $\left\{\begin{array}[]{ll}2.57^{+0.63}_{-0.50}(\beta_{B_{c}})^{+0.23}_{-0.22}(f_{M})^{+0.71}_{-0.63}(B_{K_{1}})^{+0.11}_{-0.09}(V_{cb})\times 10^{-4}&\vskip 3.41418pt\\ 4.00^{+1.01}_{-0.78}(\beta_{B_{c}})^{+0.40}_{-0.38}(f_{M})^{+0.58}_{-0.54}(B_{K_{1}})^{+0.16}_{-0.14}(V_{cb})\times 10^{-4}&\end{array}\right$., (60)
${\rm BR}(B_{c}^{+}\to\eta_{c}K_{1}(1400)^{+})$ $=$ $\left\{\begin{array}[]{ll}2.10^{+0.63}_{-0.45}(\beta_{B_{c}})^{+0.27}_{-0.26}(f_{M})^{+0.37}_{-0.35}(B_{K_{1}})^{+0.08}_{-0.08}(V_{cb})\times 10^{-4}&\vskip 3.41418pt\\ 6.70^{+2.48}_{-1.68}(\beta_{B_{c}})^{+0.78}_{-0.73}(f_{M})^{+2.92}_{-2.12}(B_{K_{1}})^{+0.27}_{-0.24}(V_{cb})\times 10^{-5}&\end{array}\right$., (63)
where the 1st (2nd) entry in Eqs. (60) and (63) corresponds to $\theta_{K}=33^{\circ}\ (58^{\circ})$. The similar patterns to the BRs also appear in the following observables for related modes trivially.
$\frac{{\rm BR}(B_{c}^{+}\to\eta_{c}K_{1}(1270)^{+})}{{\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})}$ $=$ $\left\{\begin{array}[]{ll}0.22^{+0.10}_{-0.08}&\vskip 3.41418pt\\ 0.34^{+0.14}_{-0.11}&\\ \end{array}\right.,\qquad\frac{{\rm BR}(B_{c}^{+}\to\eta_{c}K_{1}(1400)^{+})}{{\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})}=\left\{\begin{array}[]{ll}0.18^{+0.08}_{-0.07}&\vskip 3.41418pt\\ 0.06^{+0.04}_{-0.03}&\\ \end{array}\right.$. (68)
In the above numerical results, the dominant errors arise from the uncertainties of shape parameter $\beta_{B_{c}}$ and Gegenbauer moments in the distribution amplitudes of $B_{c}$, and $K_{1A}$ and $K_{1B}$ states, respectively. The $B_{c}^{+}\to\eta_{c}K_{1}(1270,1400)^{+}$ BRs indicate that ${\rm BR}(B_{c}^{+}\to\eta_{c}K_{1}(1270)^{+})$ is consistent well with ${\rm BR}(B_{c}^{+}\to\eta_{c}K_{1}(1400)^{+})$ at $\theta_{K}\sim 33^{\circ}$ within uncertainties, while ${\rm BR}(B_{c}^{+}\to\eta_{c}K_{1}(1270)^{+})$ is significantly larger than ${\rm BR}(B_{c}^{+}\to\eta_{c}K_{1}(1400)^{+})$ at $\theta_{K}\sim 58^{\circ}$ with a factor around 6. It means that an destructive interference between $B_{c}^{+}\to\eta_{c}K_{1A}^{+}$ and $B_{c}^{+}\to\eta_{c}K_{1B}^{+}$ occurs significantly in $B_{c}^{+}\to\eta_{c}K_{1}(1400)^{+}$ at $\theta_{K}\sim 58^{\circ}$. In other words, future measurements testing these $B_{c}^{+}\to\eta_{c}K_{1}(1270,1400)^{+}$ BRs with clearly different results would lead to a better determination of $\theta_{K}$. A verified $\theta_{K}$ value is key to further guide the theoretical predictions with good precision.
For the future probe of $B_{c}^{+}\to\eta_{c}K_{1}(1270,1400)^{+}$ at LHC experiments, by inputting the data ${\cal B}(K_{1}(1270)^{+}\to K^{+}\rho^{0}(\to\pi^{\pm}\pi^{\mp}))=0.190\pm 0.065$ and ${\cal B}(K_{1}(1400)^{+}\to K^{*0}(\to K^{\pm}\pi^{\mp})\pi^{+})=0.313\pm 0.020$ ParticleDataGroup:2024cfk , the BRs of multibody channels $B_{c}^{+}\to\eta_{c}K_{1}(1270)^{+}(\to K^{+}\pi^{\pm}\pi^{\mp})$ and $B_{c}^{+}\to\eta_{c}K_{1}(1400)^{+}(\to K^{\pm}\pi^{\mp}\pi^{+})$ via resonances $K_{1}(1270,1400)$ could be deduced in the iPQCD formalism under the narrow-width approximation,
${\rm BR}(B_{c}^{+}\to\eta_{c}K_{1}(1270)^{+}(\to K^{+}\pi^{\pm}\pi^{\mp}))$ $=$ $\left\{\begin{array}[]{ll}(4.88^{+2.50}_{-2.31})\times 10^{-5}&\vskip 3.41418pt\\ (7.60^{+3.51}_{-3.25})\times 10^{-5}&\end{array}\right$., (71)
${\rm BR}(B_{c}^{+}\to\eta_{c}K_{1}(1400)^{+}(\to K^{\pm}\pi^{\mp}\pi^{+}))$ $=$ $\left\{\begin{array}[]{ll}(6.57^{+2.48}_{-2.02})\times 10^{-5}&\vskip 3.41418pt\\ (2.10^{+1.23}_{-0.89})\times 10^{-5}&\end{array}\right$., (74)
It is worth emphasizing that, the determination of $\theta_{K}$ with definite value is highly important, because, if $\theta_{K}$ could be determined unambiguously, it could further help constrain the mixing between $f_{1}(1285)(h_{1}(1170))$ and $f_{1}(1420)(h_{1}(1450))$ promisingly Cheng:2011pb . The ratios between the $B_{c}^{+}\to\eta_{c}K_{1}(1270)^{+}$ and $B_{c}^{+}\to\eta_{c}K_{1}(1400)^{+}$ BRs are then derived to provide necessary reference for constraining the magnitude of $\theta_{K}$,
$\frac{{\rm BR}(B_{c}^{+}\to\eta_{c}K_{1}(1400)^{+})}{{\rm BR}(B_{c}^{+}\to\eta_{c}K_{1}(1270)^{+})}$ $=$ $\left\{\begin{array}[]{ll}0.82^{+0.09}_{-0.08}&\vskip 3.41418pt\\ 0.17^{+0.04}_{-0.04}&\\ \end{array}\right$., (77)
The above BRs and the associated ratios predicted in the iPQCD formalism would be helpful to explore the QCD dynamics in the considered axial-vectors, especially in $K_{1}(1270,1400)$.
III.2 Factorizable-emission suppressed $B_{c}^{+}\to\eta_{c}L^{+}$
Now, we turn to analyze the factorizable-emission suppressed $B_{c}^{+}\to\eta_{c}L^{+}$ decays, including $B_{c}^{+}\to\eta_{c}b_{1}(1235)^{+}$, $B_{c}^{+}\to\eta_{c}S^{+}$, and $B_{c}^{+}\to\eta_{c}T^{+}$, respectively. In explicit words, different to the factorizable-emission dominated $B_{c}^{+}\to\eta_{c}(P,V)^{+}$ decays, the factorizable-emission contributions in the former two kinds of channels are generally suppressed due to the tiny zeroth Gegenbauer moment $a_{0,b_{1}}^{\parallel}$ in the longitudinal leading-twist $b_{1}(1235)^{+}$ distribution amplitude Yang:2005gk and the nearly-zero vector decay constant $f_{S}$ for light scalars, respectively, while the latter one with even vanished factorizable-emission amplitudes at leading order is just because of the fact that a tensor meson cannot be produced via vector currents.
For the decay $B_{c}^{+}\to\eta_{c}b_{1}(1235)^{+}$, its iPQCD BR is read as,
${\rm BR}(B_{c}^{+}\to\eta_{c}b_{1}(1235)^{+})$ $=$ $7.88^{+2.52}_{-1.86}(\beta_{B_{c}})^{+0.80}_{-0.75}(f_{M})^{+3.07}_{-2.57}(a_{b_{1}})^{+0.32}_{-0.28}(V_{cb})\times 10^{-4}\;.$ (78)
accompanied by the following two ratios,
$\frac{{\rm BR}(B_{c}^{+}\to\eta_{c}b_{1}(1235)^{+})}{{\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})}$ $=$ $0.67^{+0.39}_{-0.31}\;,\qquad R_{b_{1}/\pi}^{\eta_{c}}\equiv\frac{{\rm BR}(B_{c}^{+}\to\eta_{c}b_{1}(1235)^{+})}{{\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})}=0.39^{+0.22}_{-0.18}\;,$ (79)
Notice that, the BRs of $B_{c}^{+}\to(J/\psi,\eta_{c})b_{1}(1235)^{+}$ in the iPQCD formalism indicate an interesting relation, i.e., ${\rm BR}(B_{c}^{+}\to\eta_{c}b_{1}(1235)^{+})\simeq{\rm BR}(B_{c}^{+}\to J/\psi b_{1}(1235)^{+})\sim{\cal O}(10^{-3})$ within uncertainties, even the latter BR containing three kinds of polarization contributions. These values could be tested in the near-future experiments to help understand the QCD dynamics in the related decays, as well as in $b_{1}(1235)$.
According to the categorization of light scalar mesons in two scenarios, the CP-averaged $B_{c}^{+}\to\eta_{c}S^{+}$ BRs in the iPQCD formalism are given as,
${\rm BR}(B_{c}^{+}\to\eta_{c}a_{0}(980)^{+})$ $=$ $1.02^{+0.44}_{-0.30}(\beta_{B_{c}})^{+0.24}_{-0.21}(f_{M})^{+0.04}_{-0.05}(B_{i})^{+0.04}_{-0.04}(V_{cb})\times 10^{-7}\;,$ (80)
${\rm BR}(B_{c}^{+}\to\eta_{c}\kappa^{+})$ $=$ $7.20^{+3.16}_{-2.07}(\beta_{B_{c}})^{+1.82}_{-1.52}(f_{M})^{+0.49}_{-0.46}(B_{i})^{+0.30}_{-0.25}(V_{cb})\times 10^{-6}\;,$ (81)
and
${\rm BR}(B_{c}^{+}\to\eta_{c}a_{0}(1450)^{+})$ $=$ $\left\{\begin{array}[]{ll}6.79^{+1.61}_{-1.39}(\beta_{B_{c}})^{+4.53}_{-1.40}(f_{M})^{+1.13}_{-0.65}(B_{i})^{+0.27}_{-0.24}(V_{cb})\times 10^{-9}&\vskip 3.41418pt\\ 2.15^{+0.83}_{-0.57}(\beta_{B_{c}})^{+1.00}_{-1.00}(f_{M})^{+0.05}_{-0.04}(B_{i})^{+0.09}_{-0.07}(V_{cb})\times 10^{-7}&\\ \end{array}\right$., (84)
${\rm BR}(B_{c}^{+}\to\eta_{c}{K_{0}^{*}}(1430)^{+})$ $=$ $\left\{\begin{array}[]{ll}6.68^{+1.70}_{-1.38}(\beta_{B_{c}})^{+3.39}_{-2.06}(f_{M})^{+0.21}_{-0.16}(B_{i})^{+0.28}_{-0.23}(V_{cb})\times 10^{-7}&\vskip 3.41418pt\\ 1.29^{+0.50}_{-0.34}(\beta_{B_{c}})^{+0.75}_{-0.47}(f_{M})^{+0.04}_{-0.04}(B_{i})^{+0.06}_{-0.04}(V_{cb})\times 10^{-5}&\\ \end{array}\right$., (87)
where the 1st (2nd) entry in Eqs. (84) and (87) corresponds to $S1(S2)$ and the dominant errors come mainly from the $B_{c}$-meson shape parameter $\beta_{B_{c}}$, and from the scalar decay constant $\bar{f}_{S}$ and the Gegenbauer moments $B_{i}(i=1,3)$ of scalar mesons, respectively. The future precise constraints from experimental measurements and Lattice QCD calculations on these hadronic parameters are urgently demanded for theoretical predictions with good precision.
In principle, the above-mentioned large uncertainties from nonperturbative inputs tend to be greatly cancelled by relative ratios between the relevant BRs. Then, the ratios between the $\Delta S=1$ and $\Delta S=0$ BRs in $B_{c}^{+}\to\eta_{c}S^{+}$ are easily given as,
$\frac{{\rm BR}(B_{c}^{+}\to\eta_{c}\kappa^{+})}{{\rm BR}(B_{c}^{+}\to\eta_{c}a_{0}(980)^{+})}$ $=$ $(0.71^{+0.02}_{-0.01})\times 10^{2}\;,\qquad\frac{{\rm BR}(B_{c}^{+}\to\eta_{c}{K_{0}^{*}}(1430)^{+})}{{\rm BR}(B_{c}^{+}\to\eta_{c}a_{0}(1450)^{+})}=\left\{\begin{array}[]{ll}(0.98^{+0.08}_{-0.16})\times 10^{2}&\vskip 3.41418pt\\ (0.60^{+0.11}_{-0.01})\times 10^{2}&\\ \end{array}\right.$. (90)
It seems that the errors in these results induced by the hadronic parameters are indeed cancelled to a great extent. However, it is surprisingly noted that, drastically different to the values presented in Eqs. (40) and (50), the above two ratios are significantly large near ${\cal O}(10^{2})$ within uncertainties, even already with a known factor of $|V_{us}/V_{ud}|^{2}\sim 0.05$. Moreover, they are also remarkably larger than those in the $B_{c}^{+}\to J/\psi S^{+}$ decays correspondingly within the iPQCD framework, for detail, see equation (61) in Liu:2023kxr .
Table 4: Decay amplitudes (in units of $10^{-3}$GeV-3) of $B_{c}^{+}\to\eta_{c}S^{+}$ from different twists in the iPQCD formalism. The upper (lower) entry corresponds to the scalars $a_{0}(1450)^{+}$ and ${K_{0}^{*}}(1430)^{+}$ in scenario 1 (2) at every line. For simplicity, only the central values are quoted for clarifications.
Modes Decay Amplitudes ($F_{e}$) Decay Amplitudes ($M_{e}$)
$\phi_{\eta_{c}}^{v}(x)$ $\phi_{\eta_{c}}^{s}(x)$ $\phi_{\eta_{c}}^{v}(x)$ $\phi_{\eta_{c}}^{s}(x)$
$B_{c}^{+}\to\eta_{c}a_{0}(980)^{+}$ $0.57-{\it i}1.65$ $0.25-{\it i}0.23$ $62+{\it i}2.80 0.0$$03-{\it i}0.04$
$B_{c}^{+}\to\eta_{c}\kappa^{+}$ $08-{\it i}14.79 5.5$$54+{\it i}24.88$ $1.76-{\it i}2.28$ $0.25-{\it i}0.35$
$B_{c}^{+}\to\eta_{c}a_{0}(1450)^{+}$ $\begin{array}[]{cc}-0.08-{\it i}0.21\\ -0.67-{\it i}1.91\end{array}$ $\begin{array}[]{cc}0.10+{\it i}0.34\\ 0.76+{\it i}3.32\end{array}$ $\begin{array}[]{cc}-0.15+{\it i}0.06\\ 0.20-{\it i}0.03\end{array}$ $\begin{array}[]{cc}0.01+{\it i}0.02\\ -0.03-{\it i}0.02\end{array}$
$B_{c}^{+}\to\eta_{c}{K_{0}^{*}}(1430)^{+}$ $\begin{array}[]{cc}-0.97-{\it i}2.59\\ -5.26-{\it i}15.07\end{array}$ $\begin{array}[]{cc}1.18+{\it i}4.19\\ 5.97+{\it i}26.08\end{array}$ $\begin{array}[]{cc}-1.17+{\it i}0.55\\ 1.63-{\it i}0.39\end{array}$ $\begin{array}[]{cc}0.13+{\it i}0.13\\ -0.22-{\it i}0.14\end{array}$
In order to understand this peculiar feature in the $B_{c}^{+}\to\eta_{c}S^{+}$ decays, the related decay amplitudes calculated in the iPQCD formalism are presented explicitly in Table 4. Evidently different to the $B_{c}^{+}\to J/\psi S^{+}$ decay amplitudes as shown in Table II of Liu:2023kxr , the originally naive anticipation of constructive interferences between the nonfactorizable-emission diagrams Fig. 1(c) and 1(d) with anti-symmetric leading-twist distribution amplitude of light scalars does not appear. In contrast, the sharply destructive interferences result in much smaller decay amplitudes as exhibited in the third column of Table 4. Furthermore, the destructions in $S1$ are heavier than those in $S2$ for the $B_{c}^{+}\to\eta_{c}(a_{0}(1450),K_{0}^{*}(1430))^{+}$ decays. Because of the allowed $SU(3)$ symmetry breaking effects, the comparably large factorizable-emission contributions, though which are really suppressed to the $B_{c}^{+}\to\eta_{c}(P,V)^{+}$ decays, lead to further significant destructions between the decay amplitudes of factorizable-emission and nonfactorizable-emission topologies in these $B_{c}^{+}\to\eta_{c}S^{+}$ decays. However, it should be stressed that, due to the light quark masses objectively satisfying the relation, that is, $m_{s}\gg m_{d}\sim m_{u}$, the evidently larger factorizable-emission contributions induced by the vector decay constants $f_{\kappa^{+}},f_{K_{0}^{*}(1430)^{+}}\propto(m_{s}-m_{u})$ Cheng:2005nb are therefore produced in $B_{c}^{+}\to\eta_{c}(\kappa,K_{0}^{*}(1430))^{+}$ channels, relative to the factorizable-emission amplitudes in $B_{c}^{+}\to\eta_{c}a_{0}(980,1450)^{+}$ modes proportional to the value of $(m_{d}-m_{u})$, namely, tiny broken isospin symmetry.
Nevertheless, the large iPQCD values of ${\rm BR}(B_{c}^{+}\to\eta_{c}\kappa^{+})$ and ${\rm BR}(B_{c}^{+}\to\eta_{c}K_{0}^{*}(1430)^{+})$ in $S2$ might be tested at the future LHC experiments via resonances $\kappa$ and $K_{0}^{*}(1430)$ with ${\cal B}(\kappa^{+}\to K^{0}\pi^{+})\sim 2/3$ and ${\cal B}(K_{0}^{*}(1430)^{+}\to K^{0}\pi^{+})\sim 0.62$ Maiani:2025hsz under the narrow-width approximation,
${\rm BR}(B_{c}^{+}\to\eta_{c}\kappa^{+}(\to K^{0}\pi^{+}))$ $\equiv$ ${\rm BR}(B_{c}^{+}\to\eta_{c}\kappa^{+})\cdot{\cal B}(\kappa^{+}\to K^{0}\pi^{+})=(4.80^{+2.46}_{-1.75})\times 10^{-6}\;,$ (91)
and
${\rm BR}(B_{c}^{+}\to\eta_{c}K_{0}^{*}(1430)^{+}(\to K^{0}\pi^{+}))$ $\equiv$ ${\rm BR}(B_{c}^{+}\to\eta_{c}K_{0}^{*}(1430)^{+})\cdot{\cal B}(K_{0}^{*}(1430)^{+}\to K^{0}\pi^{+})$ (92)
$=$ $(8.00^{+5.64}_{-3.60})\times 10^{-6}\;.$
The above iPQCD results predicted for the $B_{c}^{+}\to\eta_{c}S^{+}$ decays await the future experimental tests, which could help us to further explore the complicated QCD dynamics potentially.
As aforementioned, tensor states cannot be produced via the vector current. Therefore, in the $B_{c}^{+}\to\eta_{c}T^{+}$ modes, the factorizable-emission like contributions are forbidden at leading order naturally and their studies must go beyond naive factorization. Then, the $B_{c}^{+}\to\eta_{c}T^{+}$ BRs calculated in the iPQCD formalism are as follows,
${\rm BR}(B_{c}^{+}\to\eta_{c}a_{2}(1320)^{+})$ $=$ $1.33^{+0.41}_{-0.32}(\beta_{B_{c}})^{+0.16}_{-0.16}(f_{M})^{+0.05}_{-0.05}(V_{cb})\times 10^{-4}\;,$ (93)
${\rm BR}(B_{c}^{+}\to\eta_{c}{K_{2}^{*}}(1430)^{+})$ $=$ $8.21^{+2.56}_{-2.00}(\beta_{B_{c}})^{+0.79}_{-0.76}(f_{M})^{+0.33}_{-0.29}(V_{cb})\times 10^{-6}\;,$ (94)
associated with the ratios,
$\frac{{\rm BR}(B_{c}^{+}\to\eta_{c}a_{2}(1320)^{+})}{{\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})}$ $=$ $0.11^{+0.05}_{-0.04}\;,\qquad R^{\eta_{c}}_{a_{2}/\pi}\equiv\frac{{\rm BR}(B_{c}^{+}\to\eta_{c}a_{2}(1320)^{+})}{{\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})}=0.07^{+0.00}_{-0.01}\;.$ (95)
where the dominant errors come from the $B_{c}$-meson shape parameter $\beta_{B_{c}}$.
Based upon the strong decay rate, i.e., ${\cal B}(a_{2}(1320)^{+}\to\pi^{+}\pi^{-}\pi^{+})=0.351\pm 0.135$ ParticleDataGroup:2024cfk , a large BR of $B_{c}^{+}\to\eta_{c}a_{2}(1320)^{+}(\to\pi^{+}\pi^{-}\pi^{+})$ via resonance $a_{2}(1320)$ could be derived directly under the narrow-width approximation in iPQCD formalism,
${\rm BR}(B_{c}^{+}\to\eta_{c}a_{2}(1320)^{+}(\to\pi^{+}\pi^{-}\pi^{+}))$ $\equiv$ ${\rm BR}(B_{c}^{+}\to\eta_{c}a_{2}(1320)^{+})\cdot{\cal B}(a_{2}(1320)^{+}\to\pi^{+}\pi^{-}\pi^{+})$ (96)
$=$ $(4.67^{+2.37}_{-2.20})\times 10^{-5}\;,$
which can be confronted with the near-future examinations in the LHC experiments. The future tests about these values just from nonfactorizable-emission contributions would help us to examine the reliability of this iPQCD formalism.
Similar to $R_{K^{*}/\rho}^{\eta_{c}}$ in the $B_{c}^{+}\to\eta_{c}V^{+}$ sector, another ratio $R^{\eta_{c}}_{K_{2}/a_{2}}$ in the $B_{c}^{+}\to\eta_{c}T^{+}$ decays is also defined by utilizing the $B_{c}^{+}\to\eta_{c}a_{2}(1320)^{+}$ and $B_{c}^{+}\to\eta_{c}{K_{2}^{*}}(1430)^{+}$ BRs in the iPQCD framework, and its value is then read as,
$R^{\eta_{c}}_{K_{2}/a_{2}}$ $\equiv$ $\frac{{\rm BR}(B_{c}^{+}\to\eta_{c}{K_{2}^{*}}(1430)^{+})}{{\rm BR}(B_{c}^{+}\to\eta_{c}a_{2}(1320)^{+})}=(6.17^{+0.20}_{-0.13})\times 10^{-2}\;.$ (97)
This result induced by only nonfactorizable-emission contributions is very close to the value presented in Eq. (50), as well as that naively anticipated in factorization ansatz for factorizable-emission predominated $B_{c}^{+}\to\eta_{c}(\rho,K^{*})^{+}$ modes within errors.
Analogous to $B_{c}^{+}\to\eta_{c}b_{1}(1235)^{+}$ but different to $B_{c}^{+}\to\eta_{c}a_{0}(980,1450)^{+}$, the considerably constructive interferences due to the anti-symmetric leading-twist distribution amplitude between the two nonfactorizable-emission diagrams Fig. 1(c) and 1(d) work in the $B_{c}^{+}\to\eta_{c}a_{2}(1320)^{+}$ mode indeed. The future measurements on the iPQCD predictions of $B_{c}^{+}\to\eta_{c}L^{+}$ in this work could help test, even differentiate the reliability of the adopted approaches in the literature, which might help us to further understand the rich QCD dynamics in $B_{c}$-meson decays through pulling together disparate ideas.
IV Conclusions and Summary
In summary, we have studied the $B_{c}$-meson decays into $\eta_{c}$ plus a light charged meson in the self-consistent iPQCD formalism at leading order of strong coupling $\alpha_{s}$. The numerical results and phenomenological insights on CP-averaged BRs in association with relative ratios are presented explicitly. The BRs of factorizable-emission dominated decays $B_{c}^{+}\to\eta_{c}(\pi,\rho)^{+}$ and $B_{c}^{+}\to\eta_{c}a_{1}(1260)^{+}$ are generally larger than those $B_{c}\to J/\psi$ decays correspondingly in the iPQCD formalism, which reveal the remarkably different QCD dynamics between these two kinds of $B_{c}$-meson decays. For (near) future tests at experiments, the BRs of multibody modes of $B_{c}^{+}\to\eta_{c}(\pi,\rho)^{+}$ via $\eta_{c}$ resonance through $\eta_{c}\to p\bar{p}$ and $\eta_{c}\to\pi^{+}\pi^{-}(\pi^{+}\pi^{-},K^{+}K^{-},p\bar{p})$ are also predicted under the narrow-width approximation. The experimental search performed with the successfully upgraded LHCb detector for the BRs around ${\cal O}(10^{-6})$ and above, for example, ${\rm BR}(B_{c}^{+}\to\eta_{c}(\to p\bar{p})\pi^{+})=(2.70^{+0.74}_{-0.59})\times 10^{-6}$, ${\rm BR}(B_{c}^{+}\to\eta_{c}(\to p\bar{p})\rho^{+})=(7.14^{+2.03}_{-1.60})\times 10^{-6}$, ${\rm BR}(B_{c}^{+}\to\eta_{c}[\to 2(\pi^{+}\pi^{-})]\pi^{+})=(1.95^{+0.59}_{-0.50})\times 10^{-5}$, ${\rm BR}(B_{c}^{+}\to\eta_{c}[\to 2(\pi^{+}\pi^{-})]\rho^{+})=(5.16^{+1.62}_{-1.34})\times 10^{-5}$, $\cdots$, are expected to help probe the relevant decay channels and explore the nature of charmonium $\eta_{c}$. For the factorizable-emission suppressed decays, the almost equal BRs of $B_{c}^{+}\to(J/\psi,\eta_{c})b_{1}(1235)^{+}$ while the surprisingly smaller ones of $B_{c}^{+}\to\eta_{c}a_{0}(980,1450)^{+}$ than those of $B_{c}^{+}\to J/\psi a_{0}(980,1450)^{+}$ need further explorations in both of theory and experiment. Moreover, the predicted ratios $R_{K/\pi}^{\eta_{c}}$ and $R_{K^{*}/\rho}^{\eta_{c}}$ matching the anticipations based on naive ansatz means that the $B_{c}^{+}\to\eta_{c}(P,V)^{+}$ decays are predominated by factorizable-emission contributions. The experimental tests on the iPQCD ratios $R_{\eta_{c}/J/\psi}^{\pi}=1.74^{+0.66}_{-0.50}$ and ${\rm BR}(B_{c}^{+}\to\eta_{c}S^{+})_{\Delta S=0}/{\rm BR}(B_{c}^{+}\to\eta_{c}S^{+})_{\Delta S=1}\sim{\cal O}(0.01)$ might be urgently demanded, which could help decipher the QCD dynamics involved in the considered $B_{c}$-meson decays, as well as in the related charmonia $\eta_{c}$ and $J/\psi$ greatly.
Acknowledgements.
X.L. thanks Vanya Belyaev and Pei-Lian Li for their helpful communications. This work is supported by the National Natural Science Foundation of China under Grant No. 11875033 and by the Qing Lan Project of Jiangsu Province (No. 9212218405).